# Spatio-Temporal Representation Decoupling and Enhancement for Federated Instrument Segmentation in Surgical Videos

### v1

@@ 1 Introduction
2 Related Work
2.1 Surgical Instrumentation Segmentation
@@ 2.2 Federated Learning
3 Methodology
3.1 Overall Pipeline
3.1.1 Problem Formulation
@@
3.1.2 Overview of FedST
Our FedST provides a PFL framework to achieve surgical instrument segmentation in a decentralized manner. It consists of Representation Separation and Cooperation (refer to RSC Sec. 3.2) and Synthesis-based Explicit Representation Quantification (refer to SERQ Sec. 3.3). The overall pipeline is shown in Algorithm 1. FedST operates in two stages: In the local-side stage, RSC is performed to personalize the local model to adapt to diverse backgrounds by focusing on temporal dynamics and site-specific background adaptation. In the server-side stage, SERQ extracts the instrument representation from synthetic data to enhance the global model and synchronize the convergence of local site models. Regarding the training procedure, we first use the synthetic dataset for pre-training to initialize local site models.
@@ 3.1.2 Overview of FedST
3.2 Representation Separation and Cooperation
@@
3.2.1 Separation and Cooperation of Visual Representation
@@
where $B$ is the position bias; c is the number of channels; and $Q=\mathrm{FC}(\bm{F}_{ts})$, $K=\mathrm{FC}(\bm{\hat{c}})$, and $V=\mathrm{FC}(\bm{\hat{c}})$ are the query, key, and value vectors obtained from FC operations.
@@
3.2.2 Textual-guided Channel Selection
@@
3.3 Explicit Representation Quantification
@@ 3.3.1 Explicit Quantification
3.3.2 Synchronizing Local Convergence
@@
4 Experiments
4.1 Datasets and Training Implementation
4.1.1 Datasets
We construct a new FL benchmark for surgical instrument segmentation to evaluate our method and promote FL research in surgical data science. Details and statistics of each site are shown in Table 1. Each site in our FL setup corresponds to one full dataset. Specifically, Site A is EndoVis2017 [2], Site B is EndoVis2018 [3], and Site C is RARP [50], all from public surgical benchmarks. Site D (Hyst-YT) and Site E (Lob-YT) are additional datasets collected and annotated by us from open-source surgical videos on YouTube, following consistent part-level segmentation protocols. For Sites A to D, the training and testing splits follow the original protocol of each dataset, as shown in Table 1. Site E is not involved in the federated training process and serves solely as an out-of-federation test set to assess model generalization. To further enrich the training corpus, we also employ the synthetic surgical dataset Sisvse [22], which is built in Unity using semantic image synthesis. This dataset offers diverse and privacy-free surgical scenes and is only used on server-side.
Table 1: Dataset details and statistics for experimental evaluation.
@@ 4.1.1 Datasets
Table 3: Detailed performance metrics for Site E (Out Fed)
Method Dice(%$\uparrow$ IOU(%$\uparrow$ HD95(pix.$\downarrow$ ASSD(pix.$\downarrow$
@@
Table 4: Ablation Study of Key Method Components on Dice.
TS Prompt CS SERQ Dice (%$\uparrow$
@@
Table 5: Comparison of personalizing different layers in the self-attention networks.
@@ 4.1.3 Implementation Details
4.2 Comparison With State-of-the-Art Methods
4.2.1 Internal Validation
@@
The results are shown in Table 2. It can be found that FL methods show better performance than local train (Dice 70.98% vs. 75.49%). This indicates that the FL is necessary for the surgical instrument segmentation task. Data centralization achieves the best performance, indicating that collaborative training from multiple sites is beneficial. However, this method overlooks privacy concerns, limiting its utilization in real-world deployments. Our FedST can achieve promising performance while keeping privacy protection, which consistently outperforms other state-of-the-art FL methods across all evaluation metrics in the average values. This is attributed to the sharing and enhancement of consistent instrument representation across sites via RSC and SERQ modules, which strengthens the global model. These contributions are particularly beneficial for sites with limited data, such as Site D, even outperforming the centralized approach.
@@
4.2.2 External Validation
To evaluate the robustness of different methods, we further conduct external validation on the out-of-federation site E. We aggregate the parameters of personalized models on the server via weighted averaging and use the resulting global model for inference, which requires only a single forward pass and avoids the computational overhead of model ensembling. As shown in Table 3, other methods nearly fail to generalize, showing considerably lower performance. Instead, without any additional adaption on Site E, our FedST achieves promising performance. This underscores that by decoupling temporal information, our method can separate the consistent instrument representation from background tissue, and therefore largely enhance the generalization capability to the surgical scenario even with distinct background appearance.
@@ 4.2.2 External Validation
4.3 Ablation Studies
4.3.1 Effectiveness of Key Components
@@
We conduct ablation experiments to validate the effectiveness of different key components in the proposed method and obtain six configurations. The results are presented in Table 4: 1st row: we train the pure FedDP model as the baseline, where no RSC and SERQ are involved. 2nd row: the temporal modeling and separation step (TS) of RSC is utilized to separate the consistent representation of each site in FL. 3rd row: we include the textual prompt embedding without the channel selection network (CS), aiming to provide lightweight site-specific guidance. 4th row: we further include CS to form a more complete RSC module, enabling the model to selectively enhance informative channels for each site. 5th row: only SERQ which explicitly quantifies synthetic dataset is utilized to facilitate the convergence in FL model fusion without any temporal modeling and separation. 6th row: adding TS with SERQ to train the model. 7th row: our proposed full FedST.
@@
The results in Table 4 clearly show the importance of each component in FedST. In the baseline model (1st row), where neither RSC nor SERQ is used, the performance is suboptimal, particularly on site E (27.61%). When the temporal modeling and decoupling step of RSC is introduced (2nd row), the performance improves, indicating that this strategy can help separate site-specific representations and address background diversity more effectively. Adding only the textual prompt (3rd row) does not lead to significant improvement due to the lack of explicit feature selection. In contrast, introducing the channel selection network (4th row) significantly enhances the performance, particularly on sites D, demonstrating its ability to guide each site in attending to its most relevant feature channels. This highlights that channel selection serves as an effective personalized mechanism to amplify the representation learning for each local site. In the 5th row, introducing the SERQ module leverages the instrument representation from the synthetic dataset to enhance the global model, ensuring convergence during model fusion and stabilizing the optimization process, especially on site D with limited data. In the 6th row of Table 4, incorporating the temporal modeling leads to consistent performance improvement across all sites. Notably, the Dice score on Site E (out-of-federation) increases significantly from 28.23% to 76.78%. This gain highlights the effectiveness of temporal consistency in modeling surgical videos, where instrument appearance and context often persist across adjacent frames. Finally, our full FedST model (7th row) peaks the best performance across all sites, including external site E (76.99%), verifying that different proposed components can complement each other to attain better results.
4.3.2 Impact of Different Personalized FL settings
@@
To evaluate proposed FL setting is superior to other FL personalized settings, we set various personalized strategies for comparison in Table 5. 1st row: all parameters are shared in FL training. 2nd row: all parameters are personalized. 3rd row: query embedding layers in the decoder are personalized. 4th row: query embedding layers in both encoder and decoder are personalized. 5th row: query embedding layers in the encoder $\rho$ which discussed in Sec. 3.3 are personalized. In Table 5, Attn.q denotes query embedding layers in the attention-based module and Other denotes all parameters in the decoder except query embedding layers.
@@
4.3.3 Impact of Different Personalized Layers in PVT Backbone
Table 6: Different Personalization Layers in Backbone
Method Dice(%$\uparrow$
@@ 4.3.3 Impact of Different Personalized Layers in PVT Backbone
Table 7: Different Indicator in Channel Selection
Method Dice(%$\uparrow$
@@
To evaluate the effectiveness of different parameter personalization strategies within the PVT backbone, we compare various site-specific adaptations in Table 6. 1st row: only the LayerNorm layers are personalized, while all other parameters are shared. 2nd row: key embedding layers within the self-attention modules are personalized. 3rd row: value embedding layers are personalized. 4th row: both key and value embedding layers are personalized. 5th row: our proposed setting, where only the query embedding layers in the PVT encoder are personalized.
@@
4.3.4 Impact of Different Indicators in Channel Selection Network
To further investigate the effectiveness of our CLIP-based indicator, we compare it with several alternative designs in the channel selection network in Table 7. The detailed experimental settings for each method are as follows: 1st row: Randomly generates an embedding vector for each site without any semantic constraint. 2nd row: Generates embeddings using a Gaussian distribution centered at the one-hot value of each site. Specifically, the mean $\mu$ is determined by the site ID , and the standard deviation $\sigma$ is set to 0.1. This approach introduces slight variations around the one-hot indicator while maintaining a loose association with the site identity. The embedding is calculated as: $\exp\left(-\frac{(x-\mu)^{2}}{2\cdot\sigma^{2}}\right)$ where $x$ represents the embedding dimensions. 3rd row: Uses a fixed one-hot encoding to represent each site. This setting assumes complete independence and orthogonality among sites. 4th row: Utilizes the CLIP-based indicator to capture the semantic relationships among different sites. This enables more informed and meaningful channel selection based on rich prior knowledge learned from CLIP.
The results in Table 7 demonstrate clear performance differences among the evaluated methods. The Random indicator fails to provide any meaningful site-specific information, leading to poor performance due to the lack of explicit guidance for channel selection (e.g., 80.83% average Dice and only 73.28% on unseen Site E). The Gaussian indicator introduces slight variations around one-hot vectors but relies on artificial site ID mappings that do not capture real semantic relationships, resulting in limited improvements (average Dice 81.14% and 73.71% on Site E). The One-hot indicator performs slightly better by explicitly differentiating sites but still assumes strict independence and orthogonality, which prevents it from modeling shared characteristics across sites. This limitation is particularly evident on unseen Site E, where the performance remains low at 73.43%. In contrast, our CLIP-based indicator effectively captures rich semantic relationships among sites by leveraging prior surgical domain knowledge, leading to consistently better results across all sites.
4.3.5 Effectiveness of SERQ in Using Synthetic Dataset
@@
4.3.6 Comparison of State-of-arts Methods Exploiting Synthetic Data for Training
@@
4.3.7 Impact of Communication Frequency
@@ 4.4 Extensive Analysis
4.4.1 Generalization on Unseen Binary Segmentation Sites
Table 8: Performance Comparison on Unseen Binary Segmentation GraSP Dataset
Method Dice(%$\uparrow$ IOU(%$\uparrow$ HD95(pix.$\downarrow$ ASSD(pix.$\downarrow$
@@ 4.4.1 Generalization on Unseen Binary Segmentation Sites
As shown in Table 8, Although binary segmentation is conceptually simpler than part segmentation, prior methods including FedAvg [13], FedRep [14], FedDP [17], FedLD [19], and FedAS [18] struggle to generalize, achieving only around 26% Dice. These results highlight the limitations of conventional FL methods when applied to unseen dataset. In contrast, our method FedST demonstrates remarkable generalization, achieving a Dice score of 90.64% without any adaption on Grasp dataset. These results again highlight the importance of temporal modeling and instrument-specific representation enhancement in surgical video analysis.
4.4.2 Efficiency Analysis
Table 9: Communication and Efficiency Comparison across Methods
@@ 4.4.2 Efficiency Analysis
We compare the communication and efficiency characteristics of our method with several representative FL baselines, as summarized in Table 9. The table reports three key metrics: (1) communication cost per round, calculated based on the number of transmitted trainable parameters; (2) training time, including both local and server-side computation; and (3) FPS, measuring the model’s inference speed on real surgical video frames.

### v2

@@ I Introduction
II Related Work
II-A Surgical Instrumentation Segmentation
@@ II-B Federated Learning
III Methodology
III-A Overall Pipeline
III-A1 Problem Formulation
@@
III-A2 Overview of FedST
Our FedST provides a PFL framework to achieve surgical instrument segmentation in a decentralized manner. It consists of Representation Separation and Cooperation (refer to RSC Sec. III-B) and Synthesis-based Explicit Representation Quantification (refer to SERQ Sec. III-C). The overall pipeline is shown in Algorithm 1. FedST operates in two stages: In the local-side stage, RSC is performed to personalize the local model to adapt to diverse backgrounds by focusing on temporal dynamics and site-specific background adaptation. In the server-side stage, SERQ extracts the instrument representation from synthetic data to enhance the global model and synchronize the convergence of local site models. Regarding the training procedure, we first use the synthetic dataset for pre-training to initialize local site models.
@@ III-A2 Overview of FedST
III-B Representation Separation and Cooperation
@@
III-B1 Separation and Cooperation of Visual Representation
@@
[…]
@@
III-B2 Textual-guided Channel Selection
@@
III-C Explicit Representation Quantification
@@ III-C1 Explicit Quantification
III-C2 Synchronizing Local Convergence
@@
IV Experiments
IV-A Datasets and Training Implementation
IV-A1 Datasets
We construct a new FL benchmark for surgical instrument segmentation to evaluate our method and promote FL research in surgical data science. Details and statistics of each site are shown in Table I. Each site in our FL setup corresponds to one full dataset. Specifically, Site A is EndoVis2017 [2], Site B is EndoVis2018 [3], and Site C is RARP [50], all from public surgical benchmarks. Site D (Hyst-YT) and Site E (Lob-YT) are additional datasets collected and annotated by us from open-source surgical videos on YouTube, following consistent part-level segmentation protocols. For Sites A to D, the training and testing splits follow the original protocol of each dataset, as shown in Table I. Site E is not involved in the federated training process and serves solely as an out-of-federation test set to assess model generalization. To further enrich the training corpus, we also employ the synthetic surgical dataset Sisvse [22], which is built in Unity using semantic image synthesis. This dataset offers diverse and privacy-free surgical scenes and is only used on server-side.
TABLE I: Dataset details and statistics for experimental evaluation.
@@ IV-A1 Datasets
TABLE III: Detailed performance metrics for Site E (Out Fed)
[…]
@@
TABLE IV: Ablation Study of Key Method Components on Dice.
[…]
@@
TABLE V: Comparison of personalizing different layers in the self-attention networks.
@@ IV-A3 Implementation Details
IV-B Comparison With State-of-the-Art Methods
IV-B1 Internal Validation
@@
The results are shown in Table II. It can be found that FL methods show better performance than local train (Dice 70.98% vs. 75.49%). This indicates that the FL is necessary for the surgical instrument segmentation task. Data centralization achieves the best performance, indicating that collaborative training from multiple sites is beneficial. However, this method overlooks privacy concerns, limiting its utilization in real-world deployments. Our FedST can achieve promising performance while keeping privacy protection, which consistently outperforms other state-of-the-art FL methods across all evaluation metrics in the average values. This is attributed to the sharing and enhancement of consistent instrument representation across sites via RSC and SERQ modules, which strengthens the global model. These contributions are particularly beneficial for sites with limited data, such as Site D, even outperforming the centralized approach.
@@
IV-B2 External Validation
To evaluate the robustness of different methods, we further conduct external validation on the out-of-federation site E. We aggregate the parameters of personalized models on the server via weighted averaging and use the resulting global model for inference, which requires only a single forward pass and avoids the computational overhead of model ensembling. As shown in Table III, other methods nearly fail to generalize, showing considerably lower performance. Instead, without any additional adaption on Site E, our FedST achieves promising performance. This underscores that by decoupling temporal information, our method can separate the consistent instrument representation from background tissue, and therefore largely enhance the generalization capability to the surgical scenario even with distinct background appearance.
@@ IV-B2 External Validation
IV-C Ablation Studies
IV-C1 Effectiveness of Key Components
@@
We conduct ablation experiments to validate the effectiveness of different key components in the proposed method and obtain six configurations. The results are presented in Table IV: 1st row: we train the pure FedDP model as the baseline, where no RSC and SERQ are involved. 2nd row: the temporal modeling and separation step (TS) of RSC is utilized to separate the consistent representation of each site in FL. 3rd row: we include the textual prompt embedding without the channel selection network (CS), aiming to provide lightweight site-specific guidance. 4th row: we further include CS to form a more complete RSC module, enabling the model to selectively enhance informative channels for each site. 5th row: only SERQ which explicitly quantifies synthetic dataset is utilized to facilitate the convergence in FL model fusion without any temporal modeling and separation. 6th row: adding TS with SERQ to train the model. 7th row: our proposed full FedST.
@@
The results in Table IV clearly show the importance of each component in FedST. In the baseline model (1st row), where neither RSC nor SERQ is used, the performance is suboptimal, particularly on site E (27.61%). When the temporal modeling and decoupling step of RSC is introduced (2nd row), the performance improves, indicating that this strategy can help separate site-specific representations and address background diversity more effectively. Adding only the textual prompt (3rd row) does not lead to significant improvement due to the lack of explicit feature selection. In contrast, introducing the channel selection network (4th row) significantly enhances the performance, particularly on sites D, demonstrating its ability to guide each site in attending to its most relevant feature channels. This highlights that channel selection serves as an effective personalized mechanism to amplify the representation learning for each local site. In the 5th row, introducing the SERQ module leverages the instrument representation from the synthetic dataset to enhance the global model, ensuring convergence during model fusion and stabilizing the optimization process, especially on site D with limited data. In the 6th row of Table IV, incorporating the temporal modeling leads to consistent performance improvement across all sites. Notably, the Dice score on Site E (out-of-federation) increases significantly from 28.23% to 76.78%. This gain highlights the effectiveness of temporal consistency in modeling surgical videos, where instrument appearance and context often persist across adjacent frames. Finally, our full FedST model (7th row) peaks the best performance across all sites, including external site E (76.99%), verifying that different proposed components can complement each other to attain better results.
IV-C2 Impact of Different Personalized FL settings
@@
To evaluate proposed FL setting is superior to other FL personalized settings, we set various personalized strategies for comparison in Table V. 1st row: all parameters are shared in FL training. 2nd row: all parameters are personalized. 3rd row: query embedding layers in the decoder are personalized. 4th row: query embedding layers in both encoder and decoder are personalized. 5th row: query embedding layers in the encoder $\rho$ which discussed in Sec. III-C are personalized. In Table V, Attn.q denotes query embedding layers in the attention-based module and Other denotes all parameters in the decoder except query embedding layers.
@@
IV-C3 Impact of Different Personalized Layers in PVT Backbone
TABLE VI: Different Personalization Layers in Backbone
[…]
@@ IV-C3 Impact of Different Personalized Layers in PVT Backbone
TABLE VII: Different Indicator in Channel Selection
[…]
@@
To evaluate the effectiveness of different parameter personalization strategies within the PVT backbone, we compare various site-specific adaptations in Table VI. 1st row: only the LayerNorm layers are personalized, while all other parameters are shared. 2nd row: key embedding layers within the self-attention modules are personalized. 3rd row: value embedding layers are personalized. 4th row: both key and value embedding layers are personalized. 5th row: our proposed setting, where only the query embedding layers in the PVT encoder are personalized.
@@
IV-C4 Impact of Different Indicators in Channel Selection Network
To further investigate the effectiveness of our CLIP-based indicator, we compare it with several alternative designs in the channel selection network in Table VII. The detailed experimental settings for each method are as follows: 1st row: Randomly generates an embedding vector for each site without any semantic constraint. 2nd row: Generates embeddings using a Gaussian distribution centered at the one-hot value of each site. Specifically, the mean $\mu$ is determined by the site ID , and the standard deviation $\sigma$ is set to 0.1. This approach introduces slight variations around the one-hot indicator while maintaining a loose association with the site identity. The embedding is calculated as: $\exp\left(-\frac{(x-\mu)^{2}}{2\cdot\sigma^{2}}\right)$ where $x$ represents the embedding dimensions. 3rd row: Uses a fixed one-hot encoding to represent each site. This setting assumes complete independence and orthogonality among sites. 4th row: Utilizes the CLIP-based indicator to capture the semantic relationships among different sites. This enables more informed and meaningful channel selection based on rich prior knowledge learned from CLIP.
The results in Table VII demonstrate clear performance differences among the evaluated methods. The Random indicator fails to provide any meaningful site-specific information, leading to poor performance due to the lack of explicit guidance for channel selection (e.g., 80.83% average Dice and only 73.28% on unseen Site E). The Gaussian indicator introduces slight variations around one-hot vectors but relies on artificial site ID mappings that do not capture real semantic relationships, resulting in limited improvements (average Dice 81.14% and 73.71% on Site E). The One-hot indicator performs slightly better by explicitly differentiating sites but still assumes strict independence and orthogonality, which prevents it from modeling shared characteristics across sites. This limitation is particularly evident on unseen Site E, where the performance remains low at 73.43%. In contrast, our CLIP-based indicator effectively captures rich semantic relationships among sites by leveraging prior surgical domain knowledge, leading to consistently better results across all sites.
IV-C5 Effectiveness of SERQ in Using Synthetic Dataset
@@
IV-C6 Comparison of State-of-arts Methods Exploiting Synthetic Data for Training
@@
IV-C7 Impact of Communication Frequency
@@ IV-D Extensive Analysis
IV-D1 Generalization on Unseen Binary Segmentation Sites
TABLE VIII: Performance Comparison on Unseen Binary Segmentation GraSP Dataset
[…]
@@ IV-D1 Generalization on Unseen Binary Segmentation Sites
As shown in Table VIII, Although binary segmentation is conceptually simpler than part segmentation, prior methods including FedAvg [13], FedRep [14], FedDP [17], FedLD [19], and FedAS [18] struggle to generalize, achieving only around 26% Dice. These results highlight the limitations of conventional FL methods when applied to unseen dataset. In contrast, our method FedST demonstrates remarkable generalization, achieving a Dice score of 90.64% without any adaption on Grasp dataset. These results again highlight the importance of temporal modeling and instrument-specific representation enhancement in surgical video analysis.
IV-D2 Efficiency Analysis
TABLE IX: Communication and Efficiency Comparison across Methods
@@ IV-D2 Efficiency Analysis
We compare the communication and efficiency characteristics of our method with several representative FL baselines, as summarized in Table IX. The table reports three key metrics: (1) communication cost per round, calculated based on the number of transmitted trainable parameters; (2) training time, including both local and server-side computation; and (3) FPS, measuring the model’s inference speed on real surgical video frames.